Systems Staff Meeting Notes 7/3/00

Systems Staff Meeting Notes
July 3, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Ying Tang
Announcements
  1. Replacements for Ying's old programming position were interviewed last week. Winston has requested hiring one of the candidates, but no resolution has been reached yet with personnel.
  2. Bill would like everyone's single point of failure list by the end of today.
  3. Facilities hired a new employee, Brian Word, who has replaced Jimmy Baker.
  4. Geoffrey Stewart's starts on Monday, July 10.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Server for Preservation was ordered last week. It's a 100 Gig useable Xeon 733. It is really bigger than they need, we might swap it with one we already have. The Preservation server should be online by the end of the summer.
  2. Fifteen more PDA's are on order. Five of each in the following three colors: black, ice & blue.
Liaisons
  1. There's a meeting scheduled for this week, we'll wait to see if there's any virus outbreaks over the holiday before canceling it due to ALA.
  2. MMM will not meet this week.
Networking
  1. The Documents Linux server is back in service. It was bought pre-configured. Tom Minton tried to install Red Hat 6.2 as an upgrade, but it would not take. Had to start anew. RAID was the problem; security holes are plugged in 6.2.
  2. Will tested the wireless system: no signal on 1st floor, 50% on 2nd floor, 100% on 4th floor & no signal on the 5th floor. He emailed Facilities about installing conduit on the 2nd & 4th floors. Will get more antennas.
  3. Smathers CD Conversion - has been painful. The shortcuts have not been working; they refer to an exec that's not there. Had to upgrade WinSpurs. Will have to run setup on each individual work station, changing the shortcut so that it does drive mapping. All of the public work stations will have to be reconfigured. Tom Caswell will have to help us with the shortcuts. The server is up & speaks IP & IPX. We have to keep the IPX up for Xerox. Once they change that protocol, we'll drop IPX completely.
  4. Erich had an odd security report on NT9.
  5. Wiring -much is undone in Documents. We're missing an 18 pair fiber run back to room 219. Room 229 was wide open - no security. The screws that were left for the rack in 124 did not fit. Coax is hanging out of the wall. Bill would like a list of what needs to be done and he will talk to Facilities.
Software Projects Status
  1. Hilton finished the Marc conversion for Gerald's microform data. Now Gerald just needs to get the file from FCLA. Hilton has become our new Marc expert.
  2. Winston ran the Summer B Student-Patron Updates. There was a rather large discrepancy - 19,500 records off. Expiration was set for July 5th. Only Summer B students that have never attended will have trouble. Winston is waiting to hear from the Registrar.
  3. Winston is working with Betsy Simpson on an NT machine trying to run FCLA CJK Z39.50 interface.
DP Coordinator
  1. New student has been hired for Summer B - her name is Jorga (pronounced "Georgia") Bailey.
Calendar
  1. Hilton is out today.
  2. Tomorrow is the July 4th holiday.
  3. Suzy is out today, working at home on Wednesday & in the office on Thursday & Friday.
  4. Rich is out July 5-26.
  5. Bobbie is out July 10-21.
  6. Debra out Thursday, July 13.
  7. Will out July 24-28.


Back to the Systems Homepage

Last updated July 17, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu